OpenStudy (anonymous):

Find the fifth roots of 32(cos 280° + i sin 280°)

OpenStudy (tkhunny):

Have you considered DeMoivre?

OpenStudy (anonymous):

\[Let z=32(\cos 280+\iota \sin 280)=32e ^{280\iota}=32e ^{2n \pi+280\iota} \] \[find z ^{\frac{ 1 }{ 5 }}\]